In my opinion, we ought to stop making our own drums and accept that outside suppliers offer, said Wim Niewindt, managing
Answer #1
Per Unit $          65,000 $        195,000 $        225,000
Manufactured
Direct Material 10.5 $        682,500 $     2,047,500 $     2,362,500
Direct Labor (7.5*70%) 5.25 $        341,250 $     1,023,750 $     1,181,250
Variable Overhead (1.50*70%) 1.05 $          68,250 $        204,750 $        236,250
Fixed Company Overhead $        169,000 $        169,000 $        169,000
Rent $        175,500 $        175,500 $        175,500
Supervision Cost $          58,500 $          58,500 $          58,500
Total Manufacturing Cost $     1,495,000 $     3,679,000 $     4,183,000
Lees: Purchase price from Outsider $     1,170,000 $     3,510,000 $     4,050,000
Advantage $        325,000 $        169,000 $        133,000
Per Unit 65000 195000 225000
Manufactured
Direct Material 10.5 =$B$3*C1 =$B$3*D1 =$B$3*E1
Direct Labor (7.5*70%) =7.5*70% =$B$4*C1 =$B$4*D1 =$B$4*E1
Variable Overhead (1.50*70%) =1.5*70% =$B$5*C1 =$B$5*D1 =$B$5*E1
Fixed Company Overhead =2.6*65000 =C6 =D6
Rent 175500 =C7 =D7
Supervision Cost 58500 =C8 =D8
Total Manufacturing Cost =SUM(C3:C8) =SUM(D3:D8) =SUM(E3:E8)
Lees: Purchase price from Outsider =C1*18 =D1*18 =E1*18
Advantage =C9-C10 =D9-D10 =E9-E10